OpenStudy (anonymous):

write a fraction as a percent. 1/8= %

jimthompson5910 (jim_thompson5910):

First, convert 1/8 into a decimal number to get 1/8 = 0.125 Now multiply that decimal number by 100 to get 0.125*100 = 12.5% So 1/8 = 12.5%

OpenStudy (anonymous):

yeah im pretty sure its 12.5%

OpenStudy (anonymous):

think of how to write this in a way where you already know the percent of a fraction. How could you write this as something with 4 in the denominator? well it looks likes you divide the top and bottom by 2 so 8/2 = 4 and 1/2 = .5 so you end up with .5/4 now think of this as 1/2 * 1/4 what percentage is 1/4? 1/8 is half as large as 1/4 so you half the percentage.